# Category-based Galaxy Image Generation via Diffusion Models

### v1

@@ Abstract
Conventional galaxy generation methods rely on semi-analytical models and hydrodynamic simulations, which are highly dependent on physical assumptions and parameter tuning. In contrast, data-driven generative models do not have explicit physical parameters pre-determined, and instead learn them efficiently from observational data, making them alternative solutions to galaxy generation. Among these, diffusion models outperform Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) in quality and diversity. Leveraging physical prior knowledge to these models can further enhance their capabilities. In this work, we present GalCatDiff, the first framework in astronomy to leverage both galaxy image features and astrophysical properties in the network design of diffusion models. GalCatDiff incorporates an enhanced U-Net and a novel block entitled Astro-RAB (Residual Attention Block), which dynamically combines attention mechanisms with convolution operations to ensure global consistency and local feature fidelity. Moreover, GalCatDiff uses category embeddings for class-specific galaxy generation, avoiding the high computational costs of training separate models for each category. Our experimental results demonstrate that GalCatDiff significantly outperforms existing methods in terms of the consistency of sample color and size distributions, and the generated galaxies are both visually realistic and physically consistent. This framework will enhance the reliability of galaxy simulations and can potentially serve as a data augmentor to support future galaxy classification algorithm development.
1 Introduction
@@
Galaxies exhibit a wide range of morphologies (e.g., Hubble, 1926; Sandage, 1961; van den Bergh, 1976; Kormendy & Bender, 2012; Willett et al., 2013; Buta et al., 2015; Walmsley et al., 2022, 2023). The morphology of galaxies is closely linked to various physical properties, reflecting the cumulative effects of different physical processes throughout their evolutionary history (e.g., Buta, 2013; Conselice, 2014; Cappellari, 2016). For example, spiral galaxies tend to be bluer, exhibit prominent star formation, and are rotation-dominated; in contrast, elliptical galaxies are generally redder and dominated by velocity dispersion. (e.g., Strateva et al., 2001; Baldry et al., 2004; Kelvin et al., 2014; Zhu et al., 2024). Moreover, galaxies with different morphology exhibit distinct environmental distribution preferences (e.g., the so-called morphology-density relation): elliptical galaxies are typically found in denser environments, whereas spiral galaxies tend to reside in relatively isolated regions (e.g., Dressler, 1980; van der Wel et al., 2010). Furthermore, differences in the substructures of galaxies can also influence their evolution. For example, some studies suggest that bar structures can significantly influence the gas distribution within galaxies (e.g., Roberts et al., 1979; Regan & Teuben, 2004; Cole et al., 2014; Sormani et al., 2015; Li et al., 2022) and the activity of galactic nuclei (e.g., Shlosman et al., 1989; Ho et al., 1997; Lee et al., 2012). It has also been shown that disk galaxies with different bulge components, such as classical bulges or pseudobulges, exhibit significant differences in their disk properties, implying a co-evolution between the bulge and the whole galaxy (e.g., Hu et al., 2024a, b).
Numerous studies have shown that the morphology of a galaxy can serve as a powerful tool to understand its formation and evolution (e.g., El-Badry et al., 2018; Martin et al., 2018; Jackson et al., 2020; Zeng et al., 2021, 2024). Various galaxy generation methods are required to evaluate the consistency between theoretical predictions and observational data. Moreover, a detailed and well-performed galaxy morphology simulation on a large galaxy sample can further benefit the cosmic shear calibration for weak lensing analysis in the era of Euclid (e.g., Euclid Collaboration et al., 2022, 2023, 2024).
For example, semi-analytical models can qualitatively reproduce some statistical properties of galaxy morphologies (e.g., Guo et al., 2011; Zoldan et al., 2018), but they usually rely on oversimplified physical assumptions, such as exclusively linking the formation of bulges and elliptical galaxies to mergers (e.g., Kauffmann et al., 1993; Cole et al., 2000; Somerville et al., 2001, 2008; Croton et al., 2006; Gonzalez-Perez et al., 2014). Similarly, many recent hydrodynamic simulations have also succeeded in qualitatively reproducing some statistical trends related to galaxy morphology (e.g., Correa et al., 2017; Tacchella et al., 2019; Bignone et al., 2020; Pfeffer et al., 2023). Moreover, by employing post-processing techniques such as radiative transfer modeling, it is possible to generate realistic mock galaxy images that closely resemble the actual observations (e.g., Trayford et al., 2017; Rodriguez-Gomez et al., 2019; Baes et al., 2024). However, these hydrodynamic simulations heavily depend on many plausible sub-grid physics models, require meticulous parameter tuning, and are often sensitive to parameter space coverage (e.g., Vogelsberger et al., 2020; Crain & van de Voort, 2023). It is therefore unsurprising that these methods cannot easily fill in such a simulation-to-realism gap and often introduce systematic biases (e.g., Ravanbakhsh et al., 2017; Lanusse et al., 2021; Smith et al., 2022).
@@
Deep generative models, however, can directly learn the underlying statistical features from large observational datasets without relying on explicit physical parameter settings (e.g., Kingma, 2013; Goodfellow et al., 2014; Ho et al., 2020). Such data-driven approaches accurately capture data features and leverage the capabilities of neural networks and parallel computing to achieve rapid generation (e.g., Ramos & Baso, 2019; Holdship et al., 2021; Lanusse et al., 2021; Smith et al., 2022), effectively addressing the limitations of conventional methods.
@@
In recent years, deep generative models based on Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) have already been applied to galaxy generation tasks (Ravanbakhsh et al., 2017; Fussell & Moews, 2019; Smith & Geach, 2019; Lanusse et al., 2021; Holzschuh et al., 2022). As a representative in VAE-based algorithm, Ravanbakhsh et al. (2017) explores the application of both VAE and GAN-based methods, including the use of Conditional VAE (C-VAE; Kingma et al., 2014; Sohn et al., 2015). They used half-light radius, magnitude, and redshift as conditional inputs for model training, while finding the model has generated images with blurriness issue. Compared with VAEs, GAN-based models are found to be able to generate more realistic galaxy mock images (Holzschuh et al., 2022). The limitation of GAN-based models, on the other hand, is that images generated via GAN-based models are biased toward the training set, making large-size outputs unrepresentative of true cosmological structures. Additionally, training GANs remains challenging, with stable results achieved only when limited to three photometric bands (e.g., Smith & Geach, 2019). In summary, VAEs struggle to produce high-quality images (e.g., Van Den Oord et al., 2017; Vahdat & Kautz, 2020), while GANs often suffer from training instability, mode collapse, and lack of diversity in the generated samples (e.g., Goodfellow et al., 2014; Salimans et al., 2016; Che et al., 2016).
@@
Diffusion models (e.g., Ho et al., 2020) perhaps could serve as a better backbone when developing deep learning based galaxy image generative models. This is a recently developed generative model backbone that received popularity in a variety of research fields such as protein structure prediction and precipitation nowcasting. Researchers in these fields have explored leveraging physical priors into diffusion models to achieve physically consistent generation processes, further enhancing the model’s capabilities (e.g., Abramson et al., 2024; Gao et al., 2024). Astronomical applications of such backbones are still in their early stages (e.g., Smith et al., 2022; Zhao et al., 2023; Reddy et al., 2024). Zhao et al. (2023), for instance, demonstrates that diffusion models can support the conditional generation of astrophysical images and outperform StyleGAN (Karras et al., 2020), one of the most robust GAN models in their experiments, producing more robust outputs while offering reliable parameter constraints. In realistic mock galaxy image generation, Smith et al. (2022) first trained a diffusion model and managed to generate images with visually realistic morphologies and consistent physical properties such as galaxy sizes and flux.
@@
If such an algorithm can be developed, we further ask whether it can generate realistic mock galaxy images in a given morphological class. This is largely due to the data challenge that we developers are facing. In recent years, machine learning has been widely applied to handle galaxy-related downstream tasks such as source detection (e.g., Abraham et al., 2018; Angora et al., 2020; Ćiprijanović et al., 2021; Sánchez et al., 2023), classification (e.g., Dieleman et al., 2015; Aniyan & Thorat, 2017; Cheng et al., 2020; Walmsley et al., 2022; Xu et al., 2023) and segmentation (e.g., Boucaud et al., 2020; Ferreira et al., 2022; Arcelin et al., 2021). Although observational data are abundant, the distribution of different galaxy types in real datasets remains distinctly imbalanced. This imbalance directly affects the ability of machine learning models to learn minority classes, reducing their performance in these categories (e.g., Dieleman et al., 2015; Abraham et al., 2018; Xu et al., 2023). Labelling galaxy types itself can also be time-consuming and possibly survey-dependent: a large, annotated galaxy morphology model training sample can take several years even with the support from citizen scientists (e.g., Galaxy Zoo DECaLS; Walmsley et al., 2022). A powerful, class-embedded deep generative model can serve as an efficient data augmenter, producing realistic mock images in customized class and enhancing the robustness of downstream machine learning tasks (e.g., Holzschuh et al., 2022; Zhao et al., 2023).
@@
Given this background, we carry out this proof-of-concept study by proposing a novel data-driven framework, GalCatDiff, a diffusion-based model that supports image generation conditioned on galaxy categories. By introducing category embeddings, GalCatDiff can generate images for specific galaxy types, rather than relying on indiscriminate sampling from the overall data distribution, thus avoiding the high cost of training separate generative models for each category. Additionally, we design an enhanced U-Net, which optimizes the backbone U-Net of the diffusion model, and introduce a module named Astro-Residual Attention Block (Astro-RAB). This module combines attention mechanisms with convolutional operations and dynamically adjusts their weights during network training, enabling the model to maintain overall consistency while accurately reproducing local features. Experimental results show that GalCatDiff outperforms existing methods in generating high-quality and physically consistent galaxy images, achieving significant progress.
@@
The structure of this paper is as follows. Section 2 introduces the Denoising Diffusion Probabilistic Model (DDPM; Ho et al., 2020) we use in this work as the model backbone and the network architecture of GalCatDiff. Section 3 describes the dataset construction and experimental setup. Section 4 presents the generation results and discussions. Finally, Section 5 presents the conclusions and outlook.
2 Denoising Diffusion Probabilistic Models
@@
2.1 Model Architecture
In the inverse process of diffusion models, U-Net model is commonly used for predictions (e.g., Ho et al., 2020; Dhariwal & Nichol, 2021; Croitoru et al., 2023). U-Net is a convolutional neural network (CNN)-based architecture initially designed for medical image segmentation tasks, focusing on achieving high-precision, pixel-level annotations in images (e.g., Ronneberger et al., 2015; Siddique et al., 2021).
@@ 2.1 Model Architecture
In the GalCatDiff framework, we propose an enhanced U-Net designed to generate high-quality galaxy images based on class labels, while preserving their physical properties, as shown in Fig. 1. The main improvements include wider downsampling and upsampling layers with more skip connections and the introduction of the Astro-RAB. The additional skip connections facilitate the fusion of multi-scale features, reducing overfitting and thereby improving the model’s performance. The Astro-RAB integrates a fusion window attention (Liu et al., 2021) and convolution within the conventional residual block (He et al., 2016) to preserve the physical characteristics of galaxies, as illustrated in Fig. 2.
@@
The Astro-RAB consists of a convolutional block followed by an Attention Fusion Unit and a skip connection, with this sequence repeated twice. Astro-RAB concatenates the class embedding and time embedding, maps them to the module’s output dimensions through an MLP, and adds the result to the first sequence, thereby integrating the time and class embeddings into the feature map. The Attention Fusion Unit is composed of a convolution block and a window attention block. The convolutional layer efficiently extracts spatial features through local receptive fields and weight sharing, making it particularly effective for tasks involving local structures or patterns (e.g., Krizhevsky et al., 2012; Simonyan & Zisserman, 2014; Szegedy et al., 2015). The attention mechanism, in contrast, dynamically adjusts the focus of information across the entire feature map, capturing long-range dependencies (e.g., Bahdanau, 2014; Vaswani, 2017; Oktay, 2018). By combining the strengths of convolution and attention, the model can leverage both the local feature extraction capabilities of convolutions and the global information modeling capabilities of attention mechanisms (e.g., Liu et al., 2021; Dai et al., 2021; Pan et al., 2022; Wu et al., 2021; Wang et al., 2021), enhancing its ability to capture the physical features of galaxies. The output of the Attention Fusion Unit is the weighted sum of the convolutional and attention outputs. Both the convolutional weight and the attention weight are learnable parameters. Our U-Net model learns the optimal balance of these weights across different blocks to achieve the best performance, as discussed in further detail in Section 4.3.
3 Dataset and Experimental Setup
3.1 Data Collection

### v2

@@ Abstract
Conventional galaxy image generation methods rely on semi-analytical models and hydrodynamic simulations, which are highly dependent on physical assumptions and parameter tuning. In contrast, data-driven generative models do not have explicit physical parameters pre-determined, and instead learn them efficiently from observational data, making them alternative solutions to galaxy generation. Among these, diffusion models outperform Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) in quality and diversity. Embedding generalized physical features, such as category information, further enhances their generative capabilities. In this work, we present GalCatDiff, the first framework in astronomy to leverage both galaxy image features and astrophysical properties in the network design of diffusion models. GalCatDiff incorporates an enhanced U-Net and a novel block entitled Astro-RAB (Residual Attention Block), which dynamically combines attention mechanisms with convolution operations to ensure global consistency and local feature fidelity. Moreover, GalCatDiff uses category embeddings for class-specific galaxy generation, avoiding the high computational costs of training separate models for each category. Our experimental results demonstrate that GalCatDiff significantly outperforms existing methods in terms of the consistency of sample color and size distributions, and the generated galaxies are both visually realistic and physically consistent. This framework will enhance the reliability of galaxy simulations and can potentially serve as a data augmentor to support future galaxy classification algorithm development.
I Introduction
@@
Galaxies exhibit a wide range of morphologies (e.g., Hubble, 1926; Sandage, 1961; van den Bergh, 1976; Kormendy and Bender, 2012; Willett et al., 2013b; Buta et al., 2015b; Walmsley et al., 2022, 2023b). The morphology of galaxies is closely linked to various physical properties, reflecting the cumulative effects of different physical processes throughout their evolutionary history (e.g., Buta, 2013; Conselice, 2014a; Cappellari, 2016). For example, spiral galaxies tend to be bluer, exhibit prominent star formation, and are rotation-dominated; in contrast, elliptical galaxies are generally redder and dominated by velocity dispersion (e.g., Strateva et al., 2001; Baldry et al., 2004; Kelvin et al., 2014; Zhu et al., 2024). In addition, it has been shown that disk galaxies with different bulge components, such as classical bulges or pseudobulges, exhibit significant differences in their disk properties, implying a co-evolution between the bulge and the whole galaxy (e.g., Hu et al., 2024a, b).
Treating galaxy morphology as a proxy for galaxy diversity enables further investigation of the temporal evolution using redshifts extending to the early universe (e.g., Conselice, 2014b; Huertas-Company et al., 2024; Géron et al., 2025). A recent study, for instance, analyzed model predicted galaxy morphology (Spheroids, Disks, Irregular and Bulge+disk) and redshift of 20,000 galaxies in the JWST (James Webb Space Telescope, Gardner et al., 2023) Cosmic Evolution Early Release Science survey (CEERS, Finkelstein et al., 2022, 2023), confirmed that bulge-dominated galaxies decrease with redshifts, while peculiar galaxies increased with redshifts (Huertas-Company et al., 2024). In general, detailed morphological studies, especially accurate galaxy morphology classification over a large and comprehensive sample, should provide powerful diagnostic for understanding galaxy formation and evolution.
Classifying galaxy morphologies hence becomes a fundamental and crucial task for these studies. The advent of large-scale imaging surveys now could provide millions of galaxy images (e.g., York et al., 2000; Scoville et al., 2007; Abazajian et al., 2009; Dark Energy Survey Collaboration et al., 2016). Detailed morphological classifications of these galaxies not only enable the construction of useful statistical samples (e.g., Domínguez Sánchez et al., 2018; Cheng et al., 2021; Vega-Ferrero et al., 2021), but also facilitate the discovery of rare galaxy populations, such as superthin galaxies (e.g., Bizyaev et al., 2014, 2017). However, the vast scale of these images presents a significant classification challenge. While some efforts continue through manual inspection by professional astronomers (e.g., Nair and Abraham, 2010; Baillard et al., 2011; Buta et al., 2015a) or citizen science volunteers (e.g., Galaxy Zoo, Willett et al., 2013a, 2017), the majority of work now relies on automated methods. Moreover, since billions of galaxies are expected to be discovered from ongoing and upcoming surveys, such as LSST (LSST Collaboration et al., 2019), JWST (Gardner et al., 2023), Euclid (Euclid Collaboration et al., 2025), and CSST (CSST Collaboration et al., 2025), automated classification becomes an absolute necessity, while rendering manual classification nearly impossible. Early automated methods such as Decision Trees (e.g., Owens et al., 1996) and Locally Weighted Regression (e.g., De La Calleja and Fuentes, 2004) were able to handle small datasets (e.g., Bazell and Aha, 2001), while deep learning algorithms are more suitable to handle massive datasets as they have complex architectures and are able to extract features from high-dimensional data (e.g., Dieleman et al., 2015; Walmsley et al., 2023a; Luo et al., 2025). Galaxy Zoo DESI, for instance, managed to perform detailed morphology measurements for 8.7 million galaxies in the DESI Legacy Imaging Surveys using a variant of EfficientNetB0 model trained on Galaxy Zoo DECaLS labeled data (Walmsley et al., 2023a).
Although deep learning has achieved great progress in galaxy morphology research, model training relies heavily on the availability of unbiased, high-quality labeled training data. Well-labeled data usually rely on manual identification, performed by either experts or citizen scientists. Both approaches require top-down designation from galaxy sample selection, tagging workflow, and validation procedure, which inevitably introduce human biases (e.g., Lintott et al., 2008; Bamford et al., 2009; Willett et al., 2013b). Moreover, even if these biases can be partially resolved through careful selection of model architecture and training strategies (e.g., Walmsley et al., 2020b; Jiang et al., 2023), data complexity of the less representative sample class may still be incomparable with those that dominate the total sample.
@@
Deep generative models, however, can directly learn the underlying statistical features from large observational datasets without relying on explicit physical parameter settings (e.g., Kingma, 2013; Goodfellow et al., 2014; Ho et al., 2020). Such data-driven approaches accurately capture data features and leverage the capabilities of neural networks and parallel computing to achieve rapid generation (e.g., Ramos and Baso, 2019; Holdship et al., 2021; Lanusse et al., 2021; Smith et al., 2022), effectively addressing the limitations of conventional methods.
@@
In recent years, deep generative models based on Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) have already been applied to galaxy image generation tasks (Ravanbakhsh et al., 2017; Fussell and Moews, 2019; Smith and Geach, 2019; Lanusse et al., 2021; Holzschuh et al., 2022). As a representative VAE-based algorithm, Ravanbakhsh et al. (2017) introduced a conditional VAE (C-VAE; Kingma et al., 2014; Sohn et al., 2015) for galaxy image generation, where the model is conditioned on the galaxy’s half-light radius, magnitude, and redshift. Their results demonstrated that, when measured from the generated images, the samples reproduce morphological statistics such as ellipticity without significant bias. However, the images tend to lack fine structural details, appearing relatively blurred compared to real observations. Compared with VAEs, GAN-based models are found to be able to generate more realistic galaxy mock images (Holzschuh et al., 2022). The limitation of GAN-based models, on the other hand, is that the generated images are inherently biased toward the training set, and large-size outputs become unrepresentative of true cosmological structures because rare features absent from the training data cannot be synthesized (e.g., Smith and Geach, 2019). Additionally, training GANs remains challenging, with stable results achieved only when limited to three photometric bands (e.g., Smith and Geach, 2019). In summary, VAEs struggle to produce high-quality images (e.g., Van Den Oord et al., 2017; Vahdat and Kautz, 2020), while GANs often suffer from training instability, mode collapse, and lack of diversity in the generated samples (e.g., Goodfellow et al., 2014; Salimans et al., 2016; Che et al., 2016). Nevertheless, many GAN variants have sought to address these shortcomings; for example, StyleGAN has been applied to galaxy image generation and alleviates mode collapse to some extent (e.g., Holzschuh et al., 2022).
@@
Diffusion models (e.g., Ho et al., 2020) perhaps could serve as a better backbone when developing deep learning based galaxy image generative models. This is a recently developed generative model backbone that received popularity in a variety of research fields such as protein structure prediction and precipitation nowcasting. Researchers in these fields have explored leveraging physical priors into diffusion models to achieve physically consistent generation processes, further enhancing the model’s capabilities (e.g., Abramson et al., 2024; Gao et al., 2024). In the broader machine learning community, advanced efforts such as Stable Diffusion (Rombach et al., 2022) and Denoising Diffusion GANs (Xiao et al., 2021) have been developed, which extend diffusion models by operating in latent spaces for efficiency or by integrating adversarial training to accelerate sampling. These advances demonstrate the versatility of diffusion models. However, astronomy-specific applications of these diffusion backbones remain at an early stage (e.g., Smith et al., 2022; Zhao et al., 2023; Reddy et al., 2024). Zhao et al. (2023), for instance, demonstrates that diffusion models can support the conditional generation of astrophysical images and outperform StyleGAN (Karras et al., 2020), one of the most robust GAN models in their experiments, producing more robust outputs while offering reliable parameter constraints. In realistic mock galaxy image generation, Smith et al. (2022) first trained a diffusion model and managed to generate images with visually realistic morphologies and consistent physical properties such as galaxy sizes and flux.
@@
If such an algorithm can be developed, we further ask whether it can generate realistic mock galaxy images in a given morphological class. This is largely due to the data challenge that we developers are facing. In recent years, machine learning has been widely applied to handle galaxy-related downstream tasks such as source detection (e.g., Abraham et al., 2018; Angora et al., 2020; Ćiprijanović et al., 2021; Sánchez et al., 2023), classification (e.g., Dieleman et al., 2015; Aniyan and Thorat, 2017; Cheng et al., 2020; Walmsley et al., 2022; Xu et al., 2023) and segmentation (e.g., Boucaud et al., 2020; Ferreira et al., 2022; Arcelin et al., 2021). Although observational data are abundant, the distribution of different galaxy types in real datasets remains distinctly imbalanced. This imbalance directly affects the ability of machine learning models to learn minority classes, reducing their performance in these categories (e.g., Dieleman et al., 2015; Abraham et al., 2018; Xu et al., 2023). Labelling galaxy types itself can also be time-consuming and possibly survey-dependent: a large, annotated galaxy morphology model training sample can take several years even with the support from citizen scientists (e.g., Galaxy Zoo DECaLS; Walmsley et al., 2022). A powerful, class-embedded deep generative model can serve as an efficient data augmenter, producing realistic mock images in customized class and enhancing the robustness of downstream machine learning tasks (e.g., Holzschuh et al., 2022; Zhao et al., 2023).
@@
Given this background, we carry out this proof-of-concept study by proposing a novel data-driven framework, GalCatDiff, a diffusion-based model that supports image generation conditioned on galaxy categories. By introducing category embeddings, GalCatDiff can generate images for specific galaxy types, rather than relying on indiscriminate sampling from the overall data distribution, thus avoiding the high cost of training separate generative models for each category. In standard diffusion models, the U-Net primarily consists of convolutional blocks for local feature extraction, with self-attention layers inserted at certain resolutions to complement convolutional blocks by modeling global interactions across feature maps (e.g., Ho et al., 2020). However, this design treats convolution and attention as separate blocks, without explicitly integrating their complementary strengths for joint global–local feature learning. Thus, we design an enhanced U-Net, which optimizes the backbone U-Net of the standard diffusion model, and introduce a module named Astro-Residual Attention Block (Astro-RAB). This module combines attention mechanisms with convolutional operations and dynamically adjusts their weights during network training, enabling the model to maintain overall consistency while accurately reproducing local features. Experimental results show that GalCatDiff outperforms existing methods in generating high-quality and physically consistent galaxy images, achieving significant progress.
@@
The structure of this paper is as follows. Section II introduces the Denoising Diffusion Probabilistic Model (DDPM; Ho et al., 2020) we use in this work as the model backbone and the network architecture of GalCatDiff. Section III describes the dataset construction and experimental setup. Section IV presents the generation results and discussions. Finally, Section V presents the conclusions and outlook.
II Denoising Diffusion Probabilistic Models
@@
II.1 Model Architecture
In the inverse process of diffusion models, U-Net model is commonly used for predictions (e.g., Ho et al., 2020; Dhariwal and Nichol, 2021; Croitoru et al., 2023). U-Net is a convolutional neural network (CNN)-based architecture initially designed for medical image segmentation tasks, focusing on achieving high-precision, pixel-level annotations in images (e.g., Ronneberger et al., 2015; Siddique et al., 2021).
@@ II.1 Model Architecture
In the GalCatDiff framework, we propose an enhanced U-Net designed to generate high-quality galaxy images based on class labels, while preserving their physical properties, as shown in Fig. 1. The main improvements include the introduction of the Astro-RAB, the use of multiple Astro-RABs (two or more) in each down-sampling and up-sampling stage, and additional skip connections from each convolutional or Astro-RAB layer in the down-sampling path to the corresponding layer in the up-sampling path, making the architecture wider than the standard U-Net. The additional skip connections facilitate the fusion of multi-scale features, reducing overfitting and thereby improving the model’s performance. The Astro-RAB integrates a fusion window attention (Liu et al., 2021) and convolution within the conventional residual block (He et al., 2016) to preserve the physical characteristics of galaxies, as illustrated in Fig. 2.
@@
The Astro-RAB consists of a convolutional block followed by an Attention Fusion Unit and a skip connection, with this sequence repeated twice. Astro-RAB concatenates the class embedding and time embedding, maps them to the module’s output dimensions through an MLP, and adds the result to the first sequence, thereby integrating the time and class embeddings into the feature map. The Attention Fusion Unit is composed of a convolution block and a window attention block. The convolutional layer efficiently extracts spatial features through local receptive fields and weight sharing, making it particularly effective for tasks involving local structures or patterns (e.g., Krizhevsky et al., 2012; Simonyan and Zisserman, 2014; Szegedy et al., 2015). The attention mechanism, in contrast, dynamically adjusts the focus of information across the entire feature map, capturing long-range dependencies (e.g., Bahdanau, 2014; Vaswani, 2017; Oktay, 2018). By combining the strengths of convolution and attention, the model can leverage both the local feature extraction capabilities of convolutions and the global information modeling capabilities of attention mechanisms (e.g., Liu et al., 2021; Dai et al., 2021; Pan et al., 2022; Wu et al., 2021; Wang et al., 2021), enhancing its ability to capture the physical features of galaxies. The output of the Attention Fusion Unit is the weighted sum of the convolutional and attention outputs. Both the convolutional weight and the attention weight are learnable parameters. Our U-Net model learns the optimal balance of these weights across different blocks to achieve the best performance, as discussed in further detail in Section IV.3.
III Dataset and Experimental Setup
III.1 Data Collection